# OpenFly: A Versatile Toolchain and Large-scale Benchmark for Aerial Vision-Language Navigation

### v4

@@ Abstract
Vision-Language Navigation (VLN) aims to guide agents through an environment by leveraging both language instructions and visual cues, playing a pivotal role in embodied AI. Indoor VLN has been extensively studied, whereas outdoor aerial VLN remains underexplored. The potential reason is that outdoor aerial view encompasses vast areas, making data collection more challenging, which results in a lack of benchmarks. To address this problem, we propose OpenFly, a platform comprising a versatile toolchain and large-scale benchmark for aerial VLN. Firstly, we develop a highly automated toolchain for data collection, enabling automatic point cloud acquisition, scene semantic segmentation, flight trajectory creation, and instruction generation. Secondly, based on the toolchain, we construct a large-scale aerial VLN dataset with 100k trajectories, covering diverse heights and lengths across 18 scenes. The corresponding visual data are generated using various rendering engines and advanced techniques, including Unreal Engine, GTA V, Google Earth, and 3D Gaussian Splatting (3D GS). All data exhibit high visual quality. Particularly, 3D GS supports real-to-sim rendering, further enhancing the realism of the dataset. Thirdly, we propose OpenFly-Agent, a keyframe-aware VLN model, which takes language instructions, current observations, and historical keyframes as input, and outputs flight actions directly. Extensive analyses and experiments are conducted, showcasing the superiority of our OpenFly platform and OpenFly-Agent. The toolchain, dataset, and codes will be open-sourced.
Figure 1: Overview of OpenFly. This work consists of an automatic toolchain for data generation, a large-scale aerial VLN dataset comprising 100K trajectories and instructions, and a keyframe-aware VLN model emphasizing key observations.
I Introduction
Figure 2: Framework of the automatic data generation platform. Various rendering engines and simulators are first integrated, providing diverse high-quality scenes. Built on these, several interfaces and tools are developed, enabling automated generation of trajectories and instructions.
@@ I Introduction
  • Inadequate data scale. Current datasets for aerial VLN remain relatively small, with only around 10k trajectories, which is far behind embodied manipulation datasets. For example, Open X-Embodiment [36] has collected 1M episodes of manipulation data, which has significantly accelerated the development of vision-language-action (VLA) models [38, 54, 35, 55].
To address these issues, we propose OpenFly, a comprehensive platform consisting of a versatile toolchain and a large-scale benchmark for the aerial VLN task. To enhance data diversity, the platform is established on various widely-used rendering engines and advanced techniques, i.e., UE, GTA V, Google Earth, and 3D Gaussian Splatting (3D GS). This enables us to utilize a wide range of assets as shown in Fig. 1. In particular, we use UAVs for automated patrols to capture numerous real-world images and integrate 3D GS technology into our platform, through which we reconstruct realistic 3D scenes. This empowers our platform to support a real-to-sim method, enhancing the realism of data. To improve the efficiency of data collection, we develop a versatile toolchain for automated aerial VLN data generation, as depicted in Fig. 2. Specifically, point cloud acquisition is first conducted to capture the 3D occupancy of a scene. Next, scene semantic segmentation is performed to identify and select landmarks as waypoints along the flight trajectories. Building on these tools, trajectory generation is then carried out, taking landmarks and point clouds as input, using predefined flight actions as basic units, and automatically searching for a collision-free trajectory. Finally, we feed the trajectories and corresponding UAV-egocentric images into a vision-language-model (VLM), i.e., GPT-4o, to generate linguistic instructions. The entire pipeline is highly automated, reducing the reliance on UAV pilots and annotators. To collect a large-scale dataset, we meticulously collected 18 high-quality scenes, generating various trajectories of differing heights and lengths. Benefitting from our toolchain, we are able to quickly construct a dataset of 100k samples, significantly larger than existing datasets.
Besides, we propose OpenFly-Agent, a keyframe-aware aerial VLN model trained on the OpenFly dataset. It is developed based on a VLA model OpenVLA [22] and incorporates an adaptive frame-level token-sampling mechanism to mitigate visual distraction caused by redundant historical observations. Note that this sampling mechanism is crucial for the visual encoding of aerial VLN, as UAVs fly quickly and observations change rapidly. Extensive experiments are conducted on the OpenFly dataset to evaluate numerous methods, establishing a comprehensive benchmark for the aerial VLN tasks.
Overall, our contributions can be summarized as follows:
@@
  • We have constructed a novel aerial VLN benchmark comprising 100k trajectories across 18 high-quality scenes. To the best of our knowledge, this is the largest aerial VLN benchmark to date.
@@
II Related Works
II-A Simulators for Embodied AI
In recent years, many simulators related to embodied AI have been developed, e.g., MuJoCo [46], Habitat [42], Pybullet [11], Matterport3D [6], OpenAI Gym [4], and Issac Gym [32]. They allow researchers to learn and validate algorithms in safe virtual environments, significantly advancing the development of embodied intelligence. However, most of them are designed for indoor robotic manipulation and navigation, making them unsuitable as simulators for aerial VLN tasks. Gazebo and AirSim [44] are two popular UAV simulators, while Gazebo suffers from low rendering quality. Combining AirSim and UE is a potential solution to construct aerial VLN scenes. Unfortunately, the official maintenance and updates for AirSim have been discontinued, resulting in a limited availability of compatible assets.
II-B Vision-Language Navigation Datasets
@@
Numerous datasets have been proposed to accelerate the VLN task. R2R [1] focuses on evaluating agents in unseen buildings and provides discrete navigation options. RxR [25] is a densely annotated VLN dataset, supporting multilingual instructions. TouchDown [7] and REVERIE [37] have each contributed a dataset from real-life environments, which requires a ground-based agent to follow instructions and find a target object. CVDN [45] presents a cooperative VLN dataset where agents can access the dialog history of human cooperation for inference. All the above datasets are graph-based where navigable points are predefined. LANI [33] and VLN-CE [23] propose the VLN task in continuous outdoor/indoor environments, enabling agents to move freely to any unobstructed point. Recently, a few works have tried to construct VLN datasets for aerial space. ANDH [12] establishes a dialogue-based aerial VLN dataset with bird-view images. CityNav [26] builds on the point cloud data from SensatUrban [18] and linguistic annotations from CityRefer [34], which requires a real-world 2D map to help locate specific landmarks in the instruction. AerialVLN [29] and OpenUAV [47] integrate AirSim and UE to create VLN scenes where pilots can control UAVs to generate various trajectories.
II-C Vision-Language Navigation Methods
@@
VLN methods require agents to understand and complete language instructions according to visual observations. In graph-based works [31, 49, 20, 13], agents move among predefined graph nodes. Recently, LLM-driven works [58, 57, 8] use the reasoning ability of large language models to infer the next node. Although great progress has been achieved, these methods are not applicable in real unknown environments. In contrast,  [41, 19, 24, 56] explore VLN in continuous environments for more realistic applications. More recently, a few works attempt to address the problem of aerial VLN based on UAVs, among which AerialVLN [29] proposes a lookahead guidance method to generate more reasonable ground-truth trajectories during training, and STMR [14] designs a matrix representation to improve the spatial reasoning of LLMs.
III OpenFly Data Generation Platform

### v5

# OpenFly: A Comprehensive Platform for Aerial Vision-Language Navigation

@@ Abstract
Vision-Language Navigation (VLN) aims to guide agents by leveraging language instructions and visual cues, playing a pivotal role in embodied AI. Indoor VLN has been extensively studied, whereas outdoor aerial VLN remains underexplored. The potential reason is that outdoor aerial view encompasses vast areas, making data collection more challenging, which results in a lack of benchmarks. To address this problem, we propose OpenFly, a platform comprising various rendering engines, a versatile toolchain, and a large-scale benchmark for aerial VLN. Firstly, we integrate diverse rendering engines and advanced techniques for environment simulation, including Unreal Engine, GTA V, Google Earth, and 3D Gaussian Splatting (3D GS). Particularly, 3D GS supports real-to-sim rendering, further enhancing the realism of our environments. Secondly, we develop a highly automated toolchain for aerial VLN data collection, streamlining point cloud acquisition, scene semantic segmentation, flight trajectory creation, and instruction generation. Thirdly, based on the toolchain, we construct a large-scale aerial VLN dataset with 100k trajectories, covering diverse heights and lengths across 18 scenes. Moreover, we propose OpenFly-Agent, a keyframe-aware VLN model emphasizing key observations during flight. For benchmarking, extensive experiments and analyses are conducted, evaluating several recent VLN methods and showcasing the superiority of our OpenFly platform and agent. The toolchain, dataset, and codes will be open-sourced.
Figure 1: Overview of OpenFly. This work consists of the integration of 4 rendering engines, an automatic toolchain for data generation, a large-scale aerial VLN dataset comprising 100K trajectories and instructions, and a keyframe-aware VLN model emphasizing key observations.
Keywords: Aerial Vision-Language Navigation (VLN), benchmark, toolchain
1 Introduction
Figure 2: Framework of the automatic data generation. Various rendering engines are first integrated, providing diverse high-quality scenes. Built on these, several interfaces and tools are developed, enabling automated generation of trajectories and instructions.
@@ 1 Introduction
  • Inadequate data scale. Current datasets for aerial VLN remain relatively small, with only around 10k trajectories, which is far behind embodied manipulation datasets. For example, Open X-Embodiment [19] has collected 1M episodes of manipulation data, which has significantly accelerated the development of vision-language-action (VLA) models.
To address these issues, we propose OpenFly, a comprehensive platform consisting of diverse rendering engines, a versatile toolchain, and a large-scale benchmark for the aerial VLN task. To enhance data diversity, the platform is established on various widely-used rendering engines and advanced techniques, i.e., UE, GTA V, Google Earth, and 3D Gaussian Splatting (3D GS), enabling us to utilize a wide range of assets as shown in Fig. 1. In particular, we use UAVs to capture numerous real-world images and integrate 3D GS technology into our platform to reconstruct realistic 3D scenes, empowering real-to-sim simulation. To improve the efficiency of data collection, we develop a versatile toolchain for automated aerial VLN data generation as depicted in Fig. 2. Specifically, point cloud acquisition is first conducted to capture the 3D occupancy of a scene. Next, scene semantic segmentation is performed to identify and select landmarks as waypoints along the flight trajectories. Building on these tools, trajectory generation is then carried out, taking landmarks and point clouds as input, using predefined flight actions as basic units, and automatically searching for a collision-free trajectory. Finally, we feed the trajectories and corresponding UAV-egocentric images into a vision-language-model (VLM), i.e., GPT-4o, to generate linguistic instructions. The entire pipeline is highly automated, reducing the reliance on UAV pilots and annotators. To collect a large-scale dataset, we meticulously collected 18 high-quality scenes, generating various trajectories of differing heights and lengths. Benefitting from our toolchain, we are able to quickly construct a dataset of 100k samples, significantly larger than existing datasets.
Besides, we propose OpenFly-Agent, a keyframe-aware aerial VLN model incorporating an adaptive frame-level sampling mechanism to emphasize critical observations. Note that this keyframe sampling mechanism is crucial for the visual encoding of aerial VLN, as UAVs fly quickly and observations change rapidly. Extensive experiments are conducted on the OpenFly dataset to evaluate numerous methods, establishing a comprehensive benchmark for the aerial VLN tasks. Overall, our contributions can be summarized as follows:
@@
  • We have constructed a novel aerial VLN benchmark comprising 100k trajectories across 18 high-quality scenes. To the best of our knowledge, this is the largest aerial VLN benchmark to date, and users can collect more customized data using the OpenFly platform.
@@
2 Related Works
2.1 Vision-Language Navigation Datasets
@@
Numerous datasets have been proposed to accelerate the VLN task. R2R [3] focuses on evaluating agents in unseen buildings and provides discrete navigation options. RxR [4] provides a more densely annotated VLN dataset. TouchDown [1] and REVERIE [2] have each contributed a dataset from real-life environments, which requires a ground-based agent to follow instructions and find a target object. CVDN [5] presents a cooperative VLN dataset where agents can access the history of human cooperation for inference. All the above datasets are graph-based, where navigable points are predefined. LANI [7] and VLN-CE [6] propose the VLN task in continuous outdoor/indoor environments, enabling agents to move freely to any unobstructed point. Recently, a few works have tried to construct VLN datasets for aerial space. ANDH [20] establishes a dialogue-based aerial VLN dataset with bird-view images. CityNav [21] builds on the point cloud data from SensatUrban [22] and linguistic annotations from CityRefer [23], which requires a real-world 2D map to help locate specific landmarks in the instruction. AerialVLN [17] and OpenUAV [18] integrate AirSim and UE to create VLN scenes where pilots can control UAVs to generate various trajectories.
2.2 Vision-Language Navigation Methods
@@
VLN methods require agents to understand and complete language instructions according to visual observations. In graph-based works [24, 25, 26, 27], agents move among predefined graph nodes. Recently, LLM-driven works [28, 29, 30] use the reasoning ability of large language models to infer the next node. Although great progress has been achieved, these methods are not applicable in real unknown environments. In contrast,  [31, 32, 33, 16] explore VLN in continuous environments for more realistic applications. More recently, a few works attempt to address the problem of aerial VLN based on UAVs, among which AerialVLN [17] proposes a lookahead guidance method to generate more reasonable ground-truth trajectories during training. STMR [34] designs a matrix representation to improve the spatial reasoning of LLMs. OpenUAV [18] incorporates extra human assistance aligned with the ground truth (GT) trajectory to guide the navigation.
3 Automatic Data Generation
In this section, we first describe several rendering engines and data resources, then present the developed toolchain. The whole automatic data generation framework is illustrated in Fig. 2.
3.1 Rendering Engines and Data Resources
Unreal Engine. UE is a rendering engine capable of providing highly realistic virtual environments. We meticulously select 8 urban scenes, covering more than $100km^{2}$ areas. These scenes include a variety of assets such as buildings, streets, traffic lights, vehicles, and pedestrians, encompassing a wide range of architectural styles.
GTA V. GTA V is an open-world game that is frequently used by researchers due to its highly realistic environment. The game features a meticulously crafted cityscape modeled after Los Angeles, encompassing various buildings and locations such as skyscrapers, gas stations, parks, and plazas.
Google Earth. Google Earth is a virtual globe software, which builds a 3D earth model by integrating satellite imagery and aerial photographs. From this engine, we select four urban scenes covering a total area of $53.60km^{2}$, i.e., Berkeley, Osaka, Washington, D.C., and St. Louis.
3D Gaussian Splatting. 3D GS excels in realistic scene reconstruction by capturing intricate textures. It enables the creation of highly detailed, lifelike virtual environments, significantly improving the realism of digital scenes. Due to these features, we use a hierarchical 3D GS method [35] to reconstruct and render multiple real scenes. Practically, we gathered data from five campuses across three universities using UAVs. These campus scenes include various types and styles of landmarks, such as libraries, bell towers, waterways, lakes, playgrounds, construction sites, and lawns. More details of rendering engines, data resources, and examples can be found in Appendix A.
3.2 Toolchain for Automatic Data Collection
To achieve automatic data generation, we first integrate the above rendering engines and design three unified interfaces to control the agent movement and acquire sensor data (presented in Appendix C). Based on these interfaces, we further develop a toolchain, streamlining point cloud acquisition, scene semantic segmentation, trajectory creation, and instruction generation.
3D Point Cloud Acquisition. Our OpenFly provides two methods to reconstruct the point cloud map for different scenes. 1) Rasterized Sampling Reconstruction: For UE and GTA V scenes, we customize rasterized sampling points at appropriate resolutions, and then use the developed interface to obtain the local point cloud at the sampling points and stitch them for the entire scene. 2) Image-Based Sparse Reconstruction: In 3D GS, the scene reconstruction process begins with the open-source COLMAP [36] framework, which generates a sparse point cloud from input images. We directly export and use the point clouds obtained from this step.
Scene Semantic Segmentation. VLN requires meaningful landmarks as navigation targets. Thus, we offer three semantic segmentation methods to identify landmarks. 1) 3D Scene Understanding: A sequence of top-down views of the scene is captured in a rasterized format, followed by the off-the-shelf Octree-Graph [37] to extract semantic 3D instances. 2) Point Cloud Projection and Contour Extraction: We acquire the point cloud of a scene and project the voxelized point cloud onto the ground. For each instance, its contour is segmented using OpenCV, and the maximum height of its points is used as the final height. Additionally, semantic annotations are obtained by feeding the segmented instances to GPT-4o for caption. 3) Manual Annotation: When the point cloud quality of a scene is low or finer segmentation is required, OpenFly provides an interface for manually annotating instances and semantics within the point cloud. Users can choose these methods flexibly based on their requirements. The corresponding results are shown in our appendix.
Automatic Trajectory Generation. Leveraging the point cloud map and segmentation tools, OpenFly can automatically generate VLN trajectories using the following method. First, a global voxel map $M_{global}$ is constructed from the scene point cloud. Second, a landmark is randomly chosen as the target. A starting point is selected at a certain distance from the landmark, and a point close to the landmark is chosen as the endpoint. Third, A collision-free trajectory is generated using the A* [38] pathfinding algorithm based on $M_{global}$ and customized action space. By repeatedly selecting the endpoint as the new starting point, complex trajectories can be generated. Finally, utilizing OpenFly’s interface, images corresponding to the trajectory points are obtained as visual observations. A more comprehensive introduction has been included in Appendix E.
Automatic Instruction Generation. Previous research has predominantly relied on manual annotation to generate trajectory instructions, which is not only costly but also limits the scalability of datasets. To address this issue, we propose a highly automated instruction generation method based on VLMs, e.g., GPT-4o. A straightforward method would be to submit all images to VLMs to analyze the trajectory and generate instructions. However, using all images introduces considerable computational overhead and causes significant difficulties for a VLM to understand the trajectory. Additionally, we find the ‘Forward’ action usually occupies a larger proportion of a flight trajectory, with ‘Turn Left/Turn Right’ or ‘Ascend/Descend’ actions taken when encountering key landmarks.
Therefore, we split the complete trajectory into multiple sub-trajectories according to the occurrence of non-consistent actions, extracting key actions and images for processing. Notably, slight angle adjustments often occur during flight to change the direction subtly, and a slight ‘Turn’ will be merged with subsequent ‘Forward’ actions. We submit the action sequence and the last captured image of each sub-trajectory to a VLM to generate a sub-instruction of both the action and the landmark. All sub-instructions of the same trajectory are then processed by an LLM to integrate into a complete instruction. The proposed strategy significantly improves the instruction accuracy than directly inputting all trajectory images to a VLM. To further verify the data quality, we randomly select 3K samples from the entire dataset according to the data distribution in Sec. 4.2. After manually inspecting these samples, we find that they reach a high qualification rate of 91%. The problematic data involves some vague descriptions, but it is still considered acceptable by examiners. More details of instruction generation and data quality control are provided in Appendix F and G.
4 Dataset Analysis
Figure 3: Statistical analysis of the generated data. (a) Length and height distributions of trajectories. (b) Action distributions. (c) Word cloud of nouns. (d) Word cloud of verbs.
4.1 Trajectory and Instruction Analysis
Using our toolchain, we collect a dataset of 100K trajectories, which is much larger than other VLN datasets. Compared with ground-based VLN, the aerial VLN task has more motion dimensions. Therefore, we set different trajectory lengths and flight heights to obtain rich data. Fig. 3(a) exhibits the distribution of these data, with their lengths ranging from 0 to 300 meters, and the heights ranging from 0 to 150 meters. Notably, we follow the mainstream methods [6, 17] to use discrete actions, e.g., ‘Forward’ and ‘Turn left’, for trajectory generation, where the step size is 3m for ‘Forward’ action, making its proportion much higher than that of ‘Turn’ and altitude adjustments. This highly unbalanced action distribution might cause the VLN model to overfit the dominant action. To alleviate this problem, we employ three granularities of ‘Forward’ action during trajectory generation, i.e., 3m, 6m, and 9m. Fig. 3(b) presents the action distribution of our dataset. It should be noted that the OpenFly platform also supports trajectory generation with continuous waypoints.
For instruction analysis, the vocabulary size of our dataset is 15.6K, and the average length of instructions is 59. Fig. 3(c) and Fig. 3(d) illustrate the word clouds of nouns and verbs, where ‘building’, ‘windows’, and ‘skyscraper’ are the most common references, and ‘proceed’ and ‘turn’ are the mostly used verbs for VLN. Due to the space limitation, we put more details of our dataset and other ones as well as more instruction analyses in Appendix H.
4.2 Dataset Split
Similar to previous works, we divide the dataset into three splits, i.e., Train, Test Seen, Test Unseen. For the Train split, 7 scenes under the UE rendering engine account for $75.7\%$ of all data, since UE provides the largest number of scenes, where different amounts of trajectories are sampled according to the areas of scenes. The 4 scenes created by 3D GS are also the main part of the data, accounting for nearly $20\%$ of the total amount. To ensure visual quality, we only collect data from a high-altitude perspective using Google Earth, which accounts for $4.46\%$. The Test Seen data consists of 1800 trajectories uniformly sampled from 11 previously seen UE and 3D GS scenes, and the Test Unseen data comprises 1200 trajectories uniformly generated from 3 unseen scenes, i.e., UE-smallcity, 3D GS-sjtu02, and a Los Angeles-like city in GTA V. Detailed data distributions are shown in Appendix H.
5 OpenFly-Agent
In this section, we present OpenFly-Agent, an aerial VLN model that builds upon the OpenVLA [39] baseline, since OpenVLA and aerial VLN share the same pipeline, i.e., taking images and instructions as input and generating actions. Moreover, OpenVLA is trained on 1M data, having strong abilities in instruction-following and reasoning, which is critical for VLN. In contrast, our model takes a sequence of images to indicate the observation history instead of one image in the original OpenVLA. Additionally, to mitigate visual redundancy between adjacent video frames while maintaining key information, two strategies are proposed, i.e., keyframe selection and visual token merging. First, a series of candidate keyframes is selected. Then, these keyframes are merged temporally before and after the vision encoder, resulting in a compact sequence of visual tokens. Finally, the action decoder discretizes the predicted tokens to 6 action types specific to UAVs. Fig. 4 illustrates the architecture of our OpenFly-Agent, with the details of two strategies described below.
5.1 Keyframe Selection
The length of contextual visual tokens is a major challenge for VLMs when processing videos. Many open-source VLMs use uniform frame sampling [40, 41, 42] to reduce calculation, but this strategy is not suitable for aerial VLN, since it may miss frames containing key landmarks. We notice that sudden changes in the UAV’s trajectory are often caused by the observation of landmarks in both our dataset and AerialVLN [17], which can serve as cues to determine keyframes. Therefore, we adopt a heuristic method to select keyframe candidates near the change point of the UAV’s movement.
To further ensure the precision of training data, scene segmentation maps collected in Sec. 3.2 are used on selected frames to detect landmarks. Frames containing landmarks are selected as keyframes, yielding reasonably accurate results. Note that each sudden change of actions, e.g., from ‘Forward’ to ‘Turn Left’, will produce a set of keyframes. Consequently, we obtain several sets of keyframes for a long trajectory.
5.2 Visual Token Merging
Table 1: Comparison results on the test set. ‘Random’ means randomly selecting one action to execute until the ‘stop’ action is chosen. All models are retrained using our dataset.
Method test-seen test-unseen
[…]
Random 242m 0.7% 0.8% 0% 301m 0.1% 0.1% 0%
Seq2Seq[6] 291m 0.1% 3.2% 0.6% 291m 0.1% 0.9% 0.6%
CMA[6] 144m 5.7% 30.0% 5.1% 268m 0.7% 6.6% 0.7%
AerialVLN[17] 130m 6.6% 40.8% 7.0% 227m 2.3% 19.8% 2.2%
Navid[16] 142m 9.9% 24.3% 3.9% 231m 4.2% 27.0% 3.5%
NaVila[43] 140m 14.2% 41.7% 6.7% 226m 4.5% 25.1% 2.9%
OpenFly-Agent (Ours) 115m 18.5% 50.9% 12.2% 198m 5.1% 27.3% 3.5%
For each set of candidate keyframes obtained in the previous selection process, a visual encoder maps each input image to multiple visual tokens, with each token representing the information of an image patch. Considering the potential inter-frame patch redundancy, we take a strategy that similar tokens in subsequent adjacent frames are periodically merged. Specifically, we select the first frame in a keyframe set as the reference, since it usually contains the crucial observation indicating the time for action transition. Then, we densely calculate the cosine similarities between each pair of visual tokens of the reference image and the subsequent image. Next, we merge the tokens with high similarity by averaging them. The unmerged tokens in the subsequent frame will be discarded. The merging operation is iteratively performed until the entire keyframe set has been traversed. Besides, we maintain a memory bank with a capacity of $K$ images, which follows a first-in-first-out (FIFO) policy to retain the latest keyframes. Since aerial VLN requires UAVs to perform long-distance flights based on instructions, we continue to conduct token compression within each keyframe to reduce the computational burden. The compressed visual tokens are obtained through grid pooling [44]. Notably, we keep the visual tokens of the current frame uncompressed to capture the latest visual observation, as it contains the most important information for action prediction.
6 Experiments
In this section, we first introduce evaluation metrics, and then present the experimental results and ablation studies. The implementation details of the OpenFly-Agent can be found in Appendix I.
6.1 Evaluation Metrics
Four standard metrics in VLN tasks are adopted to evaluate different methods, i.e., navigation error (NE), success rate (SR), oracle success rate (OSR), and success weighted by path length (SPL). NE measures the average deviation between the UAV’s final stopping point and the ground-truth destination. SR calculates the proportion of successful tasks, where a task is considered successful if the UAV stops within 20 m of the target [17]. In OSR, if any point on the trajectory is within 20 m of the target, the task can be considered successful. SPL calculates the success rate weighted by the ratio of the ground-truth path length to the actually-executed path length.
Table 2: Ablation study on the test-seen split. ‘KS’ and ‘VTM’ denote keyframe selection and visual token merging, respectively.
[…]
OpenVLA (baseline) 196m 3.6% 12.4% 1.9%
History + VTM 215m 12.7% 39.6% 6.4%
KS 275m 6.2% 24.1% 5.6%
KS + VTM 115m 18.5% 50.9% 12.2%
6.2 Quantitative Results
We evaluate the proposed OpenFly-Agent and multiple VLN methods on the test set, with quantitative results listed in Table 1, where Seq2Seq, CMA, and AerialVLN achieve limited success rates. In contrast, Navid [16] and NaVila [43] are two most recent VLN methods, obtaining better results and demonstrating the great potential of VLMs in aerial VLN. Our OpenFly-Agent achieves the best performance for its strong VLM backbone and the proposed strategies, while aerial VLN is an emerging and challenging task, and there is still much room for improvement. The results on the test-unseen split indicate the generalization abilities of these methods. Similarly, our method achieves better performance, exhibiting a certain degree of robustness. However, all methods are significantly degraded, indicating that more powerful models are urgently needed to be developed. More qualitative results and real-world experiments are shown in Appendix J and K.
6.3 Ablation Study
Ablation studies are conducted to evaluate the contribution of the keyframe selection and visual token merging in OpenFly-Agent. Table. 2 shows the results, where OpenVLA [39] is our baseline. Using only the current frame as an observation makes OpenVLA perform poorly in the aerial VLN task. From ‘History + VTM’ we can see that historical frames significantly improve the success rate. The keyframe selection strategy further increases the SR from 12.7% to 18.5%, demonstrating the effectiveness of key observations. Besides, the comparison between ‘KS’ and ‘KS + VTM’ indicates the great effect of our visual token merging strategy.
7 Limitations and Future Work
The proposed OpenFly platform incorporates various rendering engines/techniques to provide high-quality scenes. Specifically, this is the first attempt to use 3D GS reconstructed scenes to support real-to-sim training and testing, while in the reconstruction of large-scale areas, a few visual artifacts are inevitably present. Future work will focus on exploring more effective reconstruction methods to enhance realism in large-scale scenes. Besides, the proposed OpenFly-Agent is built upon the large VLM model architecture, which is not practical for real-time deployment on UAVs. To address this, future research should focus on developing more efficient architectures and quantization techniques.
8 Conclusion
In this work, we present OpenFly, a platform designed for large-scale data collection in aerial Vision-and-Language Navigation (VLN). OpenFly integrates multiple rendering engines and provides an automatic toolchain for data generation, enabling efficient collection of diverse, high-quality aerial VLN data. The resulting large-scale dataset comprises 100k trajectories across 18 distinct scenes, spanning a wide range of altitudes and lengths, which is significantly superior than existing ones. Furthermore, we propose OpenFly-Agent, a keyframe-aware aerial navigation model capable of directly predicting flight actions based on observations and language instructions. Extensive experiments validate the effectiveness of the proposed method, and establish a comprehensive benchmark for future advancements in aerial VLN.
Appendix
Appendix A More Details of Rendering Engines and Data Resources
In this section, more information about the used rendering engines and data resources are detailed, with several high-quality examples illustrated in Fig. 5.
Unreal Engine. UE is a rendering engine capable of providing highly realistic interactive virtual environments. This platform has undergone five iterations, and each version features comprehensive and high-quality digital assets. In UE5, we meticulously select an official sample project named ‘City Sample’, which provides us with a large urban scene covering $25.3km^{2}$ and a smaller one covering $2.7km^{2}$. These scenes include a variety of assets such as buildings, streets, traffic lights, vehicles, and pedestrians. Besides, in UE4, we prepare six more high-quality scenes. Specifically, there are two large scenes showcasing the central urban areas of Shanghai and Guangzhou, covering areas of $30.88km^{2}$ and $58.56km^{2}$, respectively. The remaining four scenes are selected from AerialVLN [17]. They have smaller areas for totally about $26.64km^{2}$. These scenes encompass a wide range of architectural styles, including both Chinese and Western influences, as well as classical and modern designs. Additionally, the UE4 engine allows us to make adjustments in scene time to achieve different appearances of scenes under varying lighting conditions.
AirSim is an open-source simulator, which provides highly realistic simulated environments for UAVs and cars. We integrate the AirSim plugin into UE4 to obtain image data easily from the perspective of a UAV.
Since AirSim does not support UE5 and stopped updating in 2022, we use the UnrealCV [45] plugin as an alternative for image acquisition in UE5. To realize a highly efficient data collection in simulated scenes, we modify the UE5 project to a C++ project, integrate the UnrealCV plugin, and package executables for multiple systems like Windows and Linux.
GTA V. GTA V is an open-world game that is frequently used by computer vision researchers due to its highly realistic and dynamic virtual environment. The game features a meticulously crafted cityscape modeled after Los Angeles, encompassing various buildings and locations such as skyscrapers, gas stations, parks, and plazas, along with dynamic traffic flows and changes in lighting and shadows.
Google Earth. Google Earth is a virtual globe software, which builds a 3D earth model by integrating satellite imagery, aerial photographs, and Geographic Information System (GIS) data. From this engine, we select four urban scenes covering a total area of $53.60km^{2}$, i.e., Berkeley, primarily consisting of traditional neighborhoods; Osaka, which features a mix of skyscrapers and historic buildings; and two areas with numerous landmarks: Washington, D.C., and St. Louis.
3D Gaussian Splatting. As a highly realistic reconstruction method, hierarchical 3D GS  [35] employs a hierarchical training and display architecture, making it particularly suitable for rendering large-scale areas. Due to these features, we use this method to reconstruct and render multiple real scenes. We utilize the DJI M30T drone as the data collection device, which offers an automated oblique photography mode, enabling us to capture a large area of real-world data with minimal manpower. Practically, we gathered data from five campuses across three universities, which are East China University of Science and Technology, Northwestern Polytechnical University, and Shanghai Jiao Tong University (referred to as ECUST, NWPU, and SJTU). These campus scenes include various types and styles of landmarks, such as libraries, bell towers, waterways, lakes, playgrounds, construction sites, and lawns. The detailed information for the five campuses is presented in Table 3.